OpenStudy (anonymous):

Find the coordinates of the vertex for the parabola whose equation is f(x) = -4x2 + 8x - 4

OpenStudy (anonymous):

for your x coordinate use x=-b/2a

OpenStudy (anonymous):

then sub that to your y for the y coordinate

OpenStudy (anonymous):

lol its multiple choice my bad haha (-2,36) (-1,-16) (2,-12) (1,0)

OpenStudy (anonymous):

.f(x) = y=-4x2 + 8x - 4 , a=-4, b=8, c=-4 x=-b/2a ==-8/2(-4)=__? then sub this to y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what did you get on youR x?

OpenStudy (anonymous):

idk....just tell me the answer please. i gave you the multiple choices dude lol i really dont have time to sit here and figure it out .

OpenStudy (anonymous):

.f(x) = y=-4x2 + 8x - 4 , a=-4, b=8, c=-4 x=-b/2a x=-8/2(-4)=__? then sub this to y......... :D x=1 yes or no? sub this to f(1)=y=__ ? hmm i got into trouble of just giving answers many times ...LOL.... :D the x is there already just find y__? ...:D